OpenStudy (anonymous):

Vectors A and B have scalar product -7.00 and their vector product has magnitude 6.00. What is the angle between these vectors?

OpenStudy (shivaniits):

amm...so according to your question....A and B are vectors such that 1) their scalar product or dot product is -7 then i could write it as: \[AB \cos \theta=-7\] 2) their vector or cross product is 6 then i could say \[AB \sin \theta=6\] we could divide two equations as \[AB \sin \theta/AB \cos \theta=6/-7=-6/7\] hence further \[\tan \theta=-6/7 or \theta=\tan^{-1} (-6/7)\] as \[\tan(-\theta)=-\tan(\theta)\] hence could write \[\theta=-\tan^{-1}(6/7)\] near about -40.6012 degrees...!!